# Integrated Sensing and Communications for Pinching-Antenna Systems (PASS)

### v1

@@ Abstract
An integrated sensing and communication (ISAC) design for pinching antenna systems (PASS) is proposed, where the pinching antennas are deployed for establishing reliable line-of-sight communication and sensing links. More particularly, a separated ISAC design is proposed for the two-waveguide PASS, where one waveguide is used to emit the joint communication and sensing signals while the other waveguide is used to receive the reflected echo signals. Based on this framework, a penalty-based alternating optimization algorithm is proposed to maximize the illumination power as well as ensure the communication quality-of-service requirement. Numerical results demonstrate that 1) the proposed PASS-ISAC scheme outperforms the other baseline schemes, and 2) the considered equal power allocation model achieves a performance comparable to the optimal power allocation.
@@ I Introduction
Motivated by the above, this paper proposes a separated ISAC design for PASS. To elaborate, the BS is connected with two dielectric waveguides, where one waveguide is used to transmit the communication and sensing signals, while the other is employed to collect the reflected echo signals from the target. We aim to maximize the illumination power at the target while satisfying the quality-of-service (QoS) requirement of the communication user by optimizing the pinching beamforming offered by the mobility of pinching antennas. A penalty-based alternating optimization (AO) algorithm is proposed to handle the non-convex optimization problem, where the optimal radiation pattern is obtained under an arbitrary antenna activation configuration. Numerical results evaluate the superiority of the proposed scheme over the baseline schemes. It also reveals that the equal power model achieves a performance comparable to the optimal power allocation.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider a PASS-ISAC system, where a dual-function base station (BS) conveys with a single-antenna communicating user while sensing a point-like target. The BS is connected with two dielectric waveguides, where one waveguide is used to emit the downlink signals, and the other is employed to receive the reflected echo signals from the target. We assume each dielectric waveguide consists of $N$ pinching antennas. To support simultaneous C&S transmission with the independent pinching beamformer design, we propose a functionality-split structure for PASS, where the transmitting antenna array is partitioned into two portions. To elaborate, the former $N_{1}$ pinching antennas serve for emitting the information-bearing signals to the communication user (referred to as the communication array), and the latter $N_{2}=N-N_{1}$ antennas (referred to as the sensing array) used to sense the target. Then, the reflected echoes from the target would be collected at the receiving antennas, which are transmitted to the BS for parameter estimation.
@@ II-A ISAC Model for Pinching Antenna Networks
Consider a coherent time block of length $T$, the BS transmits the superimposed communication and sensing waveforms to the dielectric waveguide. We assume the communication channel condition and the sensing parameters remain unchanged during one coherent time block. Thus, the emitted signal at the $t$-th time slot is given by $s(t)\in\mathbb{C}$, which is assumed to be normalized and independently distributed, i.e., $\mathbb{E}\{|s(t)|^{2}\}=1$ and $\mathbb{E}\{s(t)s^{*}(\bar{t})\}=0$. On receiving $s(t)$, the dielectric waveguide radiates the signal using the pinching beamformer. The radiated signal from the pinching antenna is given by
@@
where $P_{n}$ and $\theta_{n}$ denote the power allocation and the radiation phase shift at the $n$-th pinching antenna. For ease of implementation, the equal power allocation model is assumed, i.e., $P_{1}=\cdots=P_{N}=\frac{P_{\text{max}}}{N}$ [4]. $\theta_{n}$ is defined by $2\pi\eta_{\text{eff}}\frac{\|\psi_{0}^{\text{p}}-\psi_{n}^{\text{p}}\|}{\lambda}$, where $\psi_{0}^{\text{p}}$ denotes the location of the feed point, and $\eta_{\text{eff}}$ denotes the effective refractive index of the dielectric waveguide.
@@
In this paper, we aim to utilize the communication signal to achieve simultaneous communication and target sensing. Thus, the transmitted signal can be rewritten as $\mathbf{x}(t)=[\mathbf{w}s(t),\mathbf{v}s(t)]=[\mathbf{w}s(t),\mathbf{s}(t)]$, where $\mathbf{w}s(t)$ is the communication signal and $\mathbf{v}s(t)=\mathbf{s}(t)$ represents the equivalent sensing signal. Here, we have $\mathbf{w}=\left[\sqrt{P_{1}\alpha_{\text{c}}}e^{\jmath\theta_{1}},\cdots,\sqrt{P_{N_{1}}\alpha_{\text{c}}}e^{\jmath\theta_{N_{1}}}\right]$, and $\mathbf{v}=\left[\sqrt{P_{N_{1}+1}\alpha_{\text{c}}}e^{\jmath\theta_{N_{1}+1}},\cdots,\sqrt{P_{N}\alpha_{\text{c}}}e^{\jmath\theta_{N}}\right]$.
@@ II-A1 Communication Performance Metric
$y(t)=\mathbf{h}_{[1:N_{1}],\text{c}}^{H}\mathbf{w}s(t)+\mathbf{h}_{[N_{1}+1:N],\text{c}}^{H}\mathbf{s}(t)+n(t),$ (5)
where $n(t)\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive white Gaussian noise (AWGN) at the communication user, and $\mathbf{h}_{[m:n],\xi}$ represents a subvector of $\mathbf{h}_{\xi}$ consisting of the elements with indexes from $m$ to $n$, $(m<n,\xi\in\{\text{c},\text{s}\})$. Hence, the achievable rate of the communication user is given by
$R=\log_{2}\left(1+\frac{|\mathbf{h}_{1,\text{c}}^{H}\mathbf{w}|^{2}}{\mathbf{h}_{2,\text{c}}^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{c}}^{\text{c}}+\sigma^{2}}\right),$ (6)
where $\mathbf{R}_{\text{s}}=\mathbf{v}\mathbf{v}^{H}$. For simplicity of expression, we abbreviate $\mathbf{h}_{[1:N_{1}],\text{c}}$ and $\mathbf{h}_{[N_{1}+1:N],\text{c}}$ to $\mathbf{h}_{1,\text{c}}$ and $\mathbf{h}_{2,\text{c}}$ in the following.
@@ II-A2 Sensing Performance Metric
$P(\theta_{\text{s}},r_{\text{s}})$ $=\mathbb{E}\left\{\left|\mathbf{h}_{[1:N_{1}],\text{s}}^{H}(r_{\text{s}},\phi_{\text{s}})\mathbf{w}s(t)\right|^{2}\right\}+$
$\mathbb{E}\left\{\left|\mathbf{h}_{[N_{1}+1:N],\text{s}}^{H}(r_{\text{s}},\phi_{\text{s}})\mathbf{s}(t)\right|^{2}\right\}$
$=\mathbf{h}_{1,\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{1,\text{s}}+\mathbf{h}_{2,\text{s}}^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{s}},$ (7)
where $\mathbf{h}_{[1:N_{1}],\text{s}}(r_{\text{s}},\phi_{\text{s}})=\mathbf{h}_{1,\text{s}}$ and $\mathbf{h}_{[N_{1}+1:N],\text{s}}(r_{\text{s}},\phi_{\text{s}})=\mathbf{h}_{2,\text{s}}$.
@@ II-B Problem Formulation
In this paper, we aim to maximize the illumination power $P(\theta_{\text{s}},r_{\text{s}})$ by jointly designing the pinching beamformer and antenna-activation positions, under the transmit power budget and communication QoS requirement, which is given by
$(\text{P}1)\quad$ $\max\limits_{\mathbf{h}_{1,\text{s}},\mathbf{h}_{2,\text{s}},\mathbf{h}_{1,\text{c}},\atop\mathbf{h}_{2,\text{c}},\mathbf{w},\mathbf{R}_{\text{s}}}P(\theta_{\text{s}},r_{\text{s}})$ (8a)
s.t. $|\mathbf{w}_{[n]}|^{2}=\mathbf{R}_{[n,n],\text{s}}=\frac{P_{\text{max}}}{N},$ (8b)
$\mathbf{R}\geq R_{\text{QoS}},$ (8c)
$\mathbf{R}_{\text{s}}\succeq\mathbf{0},\ \text{rank}(\mathbf{R}_{\text{s}})=1,$ (8d)
$\mathbf{h}_{[n],\text{s}}=\mathcal{H}_{\text{s}}(n),$ (8e)
$\mathbf{h}_{[n],\text{c}}=\mathcal{H}_{\text{c}}(n),$ (8f)
$|x_{n}^{\text{p}}-x_{m}^{\text{p}}|\geq\Delta x,\ n\neq m,$ (8g)
$x_{n}^{\text{p}}\in[x_{n}^{\text{p,min}},x_{n}^{\text{p,max}}],\ n\in[1,N],$ (8h)
where $\mathcal{H}_{\text{s}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}}{\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}$, $\mathcal{H}_{\text{c}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}}{\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}$, $P_{\text{max}}$ denotes the maximal transmit power at the BS, and $R_{\text{QoS}}$ denotes the minimum tolerated achievable rate of the communication user. (8b) denotes the maximal transmit power constraint; (8c) represents the QoS requirement; (8g) and (8h) limits the activation range of each pinching antenna. Notably, the problem (P1) is challenging to solve due to the quadratic objective function and the coupling between the $\{\mathbf{h}_{1,\text{s}},\mathbf{h}_{2,\text{s}},\mathbf{h}_{1,\text{c}},\mathbf{h}_{2,\text{c}}\}$ and $\{\mathbf{w},\mathbf{R}_{\text{s}}\}$.
@@ III Pinching Beamforming Optimization
In this section, we focus on the pinching beamforming design by jointly optimizing the radiation pattern and activation position of pinching antennas. Since all the channels $\{\mathbf{h}_{1,\text{s}},\mathbf{h}_{2,\text{s}},\mathbf{h}_{1,\text{c}},\mathbf{h}_{2,\text{c}}\}$ are relative to $\mathbf{x}^{\text{p}}$, which denotes the x-axis coordinate of the transmitting pinching antennas. Thus, the problem (P1) can be rewritten as
$(\text{P}2)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{w},\mathbf{R}_{\text{s}}}\quad$ $\mathbf{h}_{1,\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{1,\text{s}}+\mathbf{h}_{2,\text{s}}^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{s}}$ (9a)
s.t. $|(\mathbf{h}_{1,\text{c}})^{H}\mathbf{w}|^{2}\geq\gamma_{\text{QoS}}[(\mathbf{h}_{2,\text{c}})^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{c}}+\sigma^{2}],$ (9b)
@@
$(\text{P}3)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{W},\mathbf{R}_{\text{s}}}$ $\text{Tr}(\mathbf{H}_{1,\text{s}}\mathbf{W})+\text{Tr}(\mathbf{H}_{2,\text{s}}\mathbf{R}_{\text{s}})$ (10a)
s.t. $\mathbf{W}_{[n,n]}=\mathbf{R}_{[n,n],\text{s}}=\frac{P_{\text{max}}}{N},$ (10b)
$\text{Tr}(\mathbf{H}_{1,\text{c}}\mathbf{W})\geq\gamma_{\text{QoS}}[\text{Tr}(\mathbf{H}_{2,\text{c}}\mathbf{R}_{\text{s}})+\sigma^{2}],$ (10c)
$\mathbf{W}\succeq\mathbf{0},\ \mathbf{R}_{\text{s}}\succeq\mathbf{0},$ (10d)
$\text{rank}(\mathbf{W})=1,\ \text{rank}(\mathbf{R}_{\text{s}})=1,$ (10e)
@@
where $\mathbf{W}=\mathbf{w}\mathbf{w}^{H}$, $\mathbf{H}_{1,\text{s}}=\mathbf{h}_{1,\text{s}}\mathbf{h}_{1,\text{s}}^{H}$, $\mathbf{H}_{2,\text{s}}=\mathbf{h}_{2,\text{s}}\mathbf{h}_{2,\text{s}}^{H}$, $\mathbf{H}_{1,\text{c}}=\mathbf{h}_{1,\text{c}}\mathbf{h}_{1,\text{c}}^{H}$, and $\mathbf{H}_{2,\text{c}}=\mathbf{h}_{2,\text{c}}\mathbf{h}_{2,\text{c}}^{H}$. In order to deal with the coupled optimization variables $\{\mathbf{x}^{\text{p}}\}$ and $\{\mathbf{W},\mathbf{R}_{\text{s}}\}$, we consider the penalty-based AO framework below.
@@ III-A Radiation Pattern Design
$(\text{P}4)\ \max\limits_{\mathbf{W},\mathbf{R}_{\text{s}}}\quad$ $\text{Tr}(\mathbf{H}_{1,\text{s}}\mathbf{W})+\text{Tr}(\mathbf{H}_{2,\text{s}}\mathbf{R}_{\text{s}})$ (11a)
@@
The problem (P4) is a SDP problem with rank-one constraints, which contain $3$ feasible linear constraints and $2$ matrix variables. According to [8, Eq. (32)], it readily knows the optimal general-rank solutions of (P4) satisfy $\text{rank}(\mathbf{W})^{2}+\text{rank}(\mathbf{R}_{\text{s}})^{2}\leq 3$, which indicates that $\text{rank}(\mathbf{W})=\text{rank}(\mathbf{R}_{\text{s}})=1$. By applying the conclusion, the optimal pinching beamformer can be obtained.

### v2

@@ Abstract
An integrated sensing and communication (ISAC) design for pinching antenna systems (PASS) is proposed, where the pinching antennas are deployed for establishing reliable line-of-sight communication and sensing links. More particularly, a separated ISAC design is proposed for the two-waveguide PASS, where one waveguide is used to emit the joint communication and sensing signals while the other waveguide is used to receive the reflected echo signals. Based on this framework, a penalty-based alternating optimization algorithm is proposed to maximize the illumination power as well as ensure the communication quality-of-service requirement. Numerical results demonstrate that 1) the proposed PASS-ISAC scheme outperforms the other baseline schemes, and 2) the considered equal power allocation model achieves an upper bound performance.
@@ I Introduction
Motivated by the above, this paper proposes a separated ISAC design for PASS. To elaborate, the base station (BS) is connected with two dielectric waveguides, where one waveguide is used to transmit the communication and sensing signals, while the other is employed to collect the reflected echo signals from the target. We aim to maximize the illumination power at the target while satisfying the quality-of-service (QoS) requirement of the communication user by optimizing the pinching beamforming offered by the mobility of pinching antennas. A penalty-based alternating optimization (AO) algorithm is proposed to handle the non-convex optimization problem, where the optimal radiation pattern is obtained under an arbitrary antenna activation configuration. Numerical results evaluate the superiority of the proposed scheme over the baseline schemes. It also reveals that the equal power model achieves a performance comparable to the upper bound.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider a PASS-ISAC system, where a dual-function BS conveys with a single-antenna communicating user while sensing a point-like target. The BS is connected with two dielectric waveguides, where one waveguide is used to emit the downlink signals, and the other is employed to receive the reflected echo signals from the target. It is assumed that each dielectric waveguide consists of $N$ pinching antennas. We aim to achieve simultaneous C&S transmission in PASS using only information-bearing signals, which are emitted from the transmitting antennas. Then, the reflected echoes from the target would be collected at the receiving antennas, which are transmitted to the BS for parameter estimation.
@@ II-A ISAC Model for Pinching Antenna Networks
Consider a coherent time block of length $T$, the BS transmits the superimposed communication and sensing waveforms to the dielectric waveguide. We assume the communication channel condition and the sensing parameters remain unchanged during one coherent time block. Thus, the emitted signal at the $t$-th time slot is given by $s(t)\in\mathbb{C}$, which is assumed to be normalized and independently distributed, i.e., $\mathbb{E}\{|s(t)|^{2}\}=1$ and $\mathbb{E}\{s(t)s^{*}(\bar{t})\}=0$. On receiving $s(t)$, the dielectric waveguide radiates the signal as follows.
@@
where $P_{n}$ and $\theta_{n}$ denote the power allocation and the radiation phase shift at the $n$-th pinching antenna. For ease of implementation, the equal power allocation model is assumed, i.e., $P_{1}=\cdots=P_{N}=\frac{P_{\text{max}}}{N}$ [4]. $\theta_{n}$ is defined by $2\pi\eta_{\text{eff}}\frac{\|\psi_{0}^{\text{p}}-\psi_{n}^{\text{p}}\|}{\lambda}$, where $\psi_{0}^{\text{p}}$ denotes the location of the feed point, and $\eta_{\text{eff}}$ denotes the effective refractive index of the dielectric waveguide. According to (2)-(4), we can observe that the positions of pinching antennas have a significant impact on both the free-space path loss of wireless channels and the phase shift of radiated signals. This indicates that it is possible to establish the favorable wireless propagation and manipulate the radiated characteristics of signals by altering the positions of pinching antennas. The capability of reconfiguring signal propagation is referred as pinching beamforming.
@@
In this paper, we aim to utilize the communication signal to achieve simultaneous communication and target sensing. Thus, the transmitted signal can be rewritten as $\mathbf{x}(t)=\mathbf{w}s(t)$, where $\mathbf{w}=\left[\sqrt{P_{1}\alpha_{\text{c}}}e^{\jmath\theta_{1}},\cdots,\sqrt{P_{N}\alpha_{\text{c}}}e^{\jmath\theta_{N}}\right]$.
@@ II-A1 Communication Performance Metric
$y(t)=\mathbf{h}_{\text{c}}^{H}\mathbf{w}s(t)+n(t),$ (5)
where $n(t)\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive white Gaussian noise (AWGN) at the communication user. Hence, the achievable rate of the communication user is given by
$R=\log_{2}\left(1+\frac{|\mathbf{h}_{\text{c}}^{H}\mathbf{w}|^{2}}{\sigma^{2}}\right).$ (6)
@@ II-A2 Sensing Performance Metric
$P(\theta_{\text{s}},r_{\text{s}})$ $=\mathbb{E}\left\{\left|\mathbf{h}_{\text{s}}^{H}(r_{\text{s}},\phi_{\text{s}})\mathbf{w}s(t)\right|^{2}\right\}$
$=\mathbf{h}_{\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{\text{s}}.$ (7)
@@ II-B Problem Formulation
In this paper, we aim to maximize the illumination power $P(\theta_{\text{s}},r_{\text{s}})$ by jointly designing the pinching beamformer, under the transmit power budget and communication QoS requirement, which is given by
$(\text{P}1)\quad$ $\max\limits_{\mathbf{h}_{\text{s}},\mathbf{h}_{\text{c}},\mathbf{w}}P(\theta_{\text{s}},r_{\text{s}})$ (8a)
s.t. $|\mathbf{w}_{[n]}|^{2}=\frac{P_{\text{max}}}{N},$ (8b)
$R\geq R_{\text{QoS}},$ (8c)
$\mathbf{h}_{[n],\text{s}}=\mathcal{H}_{\text{s}}(n),$ (8d)
$\mathbf{h}_{[n],\text{c}}=\mathcal{H}_{\text{c}}(n),$ (8e)
$|x_{n}^{\text{p}}-x_{m}^{\text{p}}|\geq\Delta x,\ n\neq m,$ (8f)
$x_{n}^{\text{p}}\in[x_{n}^{\text{p,min}},x_{n}^{\text{p,max}}],\ n\in[1,N],$ (8g)
where $\mathcal{H}_{\text{s}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}}{\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}$, $\mathcal{H}_{\text{c}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}}{\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}$, $P_{\text{max}}$ denotes the maximal transmit power at the BS, and $R_{\text{QoS}}$ denotes the minimum tolerated achievable rate of the communication user. (8b) denotes the maximal transmit power constraint; (8c) represents the QoS requirement; (8f) and (8g) limits the activation range of each pinching antenna. Notably, the problem (P1) is challenging to solve due to the quadratic objective function and the coupling between the $\{\mathbf{h}_{\text{s}},\mathbf{h}_{\text{c}}\}$ and $\{\mathbf{w}\}$.
@@ III Pinching Beamforming Optimization
In this section, we focus on the pinching beamforming design by jointly optimizing the radiation pattern and activation position of pinching antennas. Since all the channels $\{\mathbf{h}_{\text{s}},\mathbf{h}_{\text{c}}\}$ are relative to $\mathbf{x}^{\text{p}}$, which denotes the x-axis coordinate of the transmitting pinching antennas. Thus, the problem (P1) can be rewritten as
$(\text{P}2)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{w}}\quad$ $\mathbf{h}_{\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{\text{s}}$ (9a)
s.t. $|\mathbf{h}_{\text{c}}^{H}\mathbf{w}|^{2}\geq\gamma_{\text{QoS}}\sigma^{2},$ (9b)
@@
$(\text{P}3)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{W}}$ $\text{Tr}(\mathbf{H}_{\text{s}}\mathbf{W})$ (10a)
s.t. $\mathbf{W}_{[n,n]}=\frac{P_{\text{max}}}{N},$ (10b)
$\text{Tr}(\mathbf{H}_{\text{c}}\mathbf{W})\geq\gamma_{\text{QoS}}\sigma^{2},$ (10c)
$\mathbf{W}\succeq\mathbf{0},$ (10d)
$\text{rank}(\mathbf{W})=1,$ (10e)
@@
where $\mathbf{W}=\mathbf{w}\mathbf{w}^{H}$, $\mathbf{H}_{\text{s}}=\mathbf{h}_{\text{s}}\mathbf{h}_{\text{s}}^{H}$, and $\mathbf{H}_{\text{c}}=\mathbf{h}_{\text{c}}\mathbf{h}_{\text{c}}^{H}$. In order to deal with the coupled optimization variables $\{\mathbf{x}^{\text{p}}\}$ and $\{\mathbf{W}\}$, we consider the penalty-based AO framework below.
@@ III-A Radiation Pattern Design
$(\text{P}4)\ \max\limits_{\mathbf{W}}\quad$ $\text{Tr}(\mathbf{H}_{\text{s}}\mathbf{W})$ (11a)
@@
The problem (P4) is a SDP problem with rank-one constraints, which contain $2$ feasible linear constraints and $2$ matrix variables. According to [8, Eq. (32)], it readily knows the optimal general-rank solutions of (P4) satisfy
$\text{rank}(\mathbf{W})\leq\sqrt{2},$ (12)
which indicates that $\text{rank}(\mathbf{W})=1$. By applying the conclusion, the optimal radiation pattern can be obtained.